OpenStudy (waheguru):

A child is on a swing. Her mom pushes her to a height of 1.5 m above the starting position. Determine the child's maximum velocity during the swing.

OpenStudy (waheguru):

For this question, I am supposed to fine the kinetic energy, not potential gravitation energy right? Because the Kinetic energy formula has velocity so I can find the max velocity?

OpenStudy (waheguru):

With the information given, it is asking me to find kinetic energu right? so I then can find the velocity from the formula E(kin) = 1/2mv^2

OpenStudy (waheguru):

mgΔd=1/2mv^2

OpenStudy (waheguru):

then mass cancels out and I can solve for velocity

OpenStudy (waheguru):

would it be possible to get the right answer by finding potential gravitational energy?

OpenStudy (experimentx):

hmm ... use conservation of energy KE = PE

OpenStudy (experimentx):

mgh = 1/2 mv^2 v = sqrt(2*g*h)

OpenStudy (waheguru):

is it always true for KE = PE ?

OpenStudy (waheguru):

what about pendulum when it reaches max height. at that point it mas max potential grav energy but no kinetic

OpenStudy (waheguru):



OpenStudy (experimentx):

yep ... if you ignore dissipative forces like friction.

OpenStudy (experimentx):

yes for the other question too.

OpenStudy (waheguru):

how?

OpenStudy (waheguru):

KE = 0 GE > 0

OpenStudy (waheguru):

PE > 0

OpenStudy (waheguru):

can you explain further?

OpenStudy (experimentx):

hmm ... you mean why it should always be true? that KE = PE ??

OpenStudy (waheguru):

Yea because on the pendulum when the mass reaches max height, there is an instaneous pause which means there is no kinetic energy but there still is potential energy

OpenStudy (waheguru):

so in that case the KE is not equal to PE

OpenStudy (experimentx):

do you know about conservation of energy?

OpenStudy (waheguru):

yea energy is never lost

OpenStudy (experimentx):

yep ...that's why!!

OpenStudy (waheguru):

ok i think i understand now

OpenStudy (waheguru):

the kinetic energy turns into potential energy but total mechanical energy is the same

OpenStudy (waheguru):

thanks for your help

OpenStudy (experimentx):

yep :D